OpenStudy (anonymous):

A business permits its custimers to pay with either a credit card or to receive a percentage discount r for paying cash. For credit card purchases, the business receives 95% of the purchase price one-half month later. At an effective annual rate of 12%, the two payment methods are equivalent. Determine r.

OpenStudy (anonymous):

If i could just get help starting, I would appreciate it. Even just the formula.

OpenStudy (anonymous):

lets try with numbers

OpenStudy (anonymous):

hm?

OpenStudy (anonymous):

also, is percentage discount the same retriceeffective rate or discount?

OpenStudy (anonymous):

actually the more i read this the more i don't understand it. but i was going to say lets imagine that you buy something for $100 then if payed by credit card the business receives $95 two weeks later, but then i am lost after that, because i don't know who is being charged 12% interest

OpenStudy (anonymous):

I KNOW!!! f-ck theories of interest.... would it be the extra 5%? HA! do you know the formula for percentage discount by any chance?

OpenStudy (anonymous):

this line At an effective annual rate of 12%, the two payment methods are equivalent makes no sense to me. does it make sense to you?

OpenStudy (anonymous):

none whatsoever.

OpenStudy (anonymous):

does it mean the business lost 12% annually on the $95 it took two weeks to get ?

OpenStudy (anonymous):

that i can compute

OpenStudy (anonymous):

r = (1+i/n)^n - 1 i just looked up the formula for the eff. annual int rate

OpenStudy (anonymous):

yeah i get that, but i don't know how it applies to this problem

OpenStudy (anonymous):

neither do i. i have multiple choice answers of 4.55, 4.85, 5.15, 5.45, and 5.75. but i don't know how to even approach this

OpenStudy (mathmate):

5.45 is correct.

OpenStudy (anonymous):

how so?

OpenStudy (mathmate):

This is how it works: Using Satellite73's approach on a purchase of $100.

OpenStudy (mathmate):

If the customer pays by credit card, the merchant will get $95 (1/24) of a year after. Effective interest rate of 12% translates to 0.4733195% per (1/24) of a year, using your formula. If the customer pays cash, merchant get X dollars right away, which becomes $95 (1/24) of a year after.

OpenStudy (anonymous):

ok i think i see it. the money you lose is what you would have got if you invest the 100 at .12 effective for two weeks right?

OpenStudy (mathmate):

So X*1.004733195=95, or X=$94.55246 Which means that the merchant would have given a discount of 5.45%

OpenStudy (anonymous):

it is \(\frac{1}{26}\) of the year later, but i think the calculation is correct

OpenStudy (mathmate):

It said half a month later, so I used 1/24. It would be very close anyway.

OpenStudy (anonymous):

sorry, where did the (1/24) come in? is it 24 payments in the year?

OpenStudy (anonymous):

oh silly me yes, you are right for some reason i used 2 weeks i am wrong

OpenStudy (mathmate):

The question says merchant gets money 1/2 a month later. So for calcuations, we have to assume 24 periods.

OpenStudy (anonymous):

ok, i am beginning to see

OpenStudy (anonymous):

lot of assumptions here

OpenStudy (anonymous):

yea

OpenStudy (mathmate):

That's the world of business! :)

OpenStudy (anonymous):

we are also making the assumption that "the two payment methods are equivalent" means after one year they are equivalent

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

@mathmate i have a different method, let me know if it is correct

OpenStudy (mathmate):

I guess the merchant has to look at the long term. In any case, 1/2 month could easily mean anything form 13 to 17 days, when we count the weekends. @IStutts you're welcome. Let's look at Satellite73's alternative method.

OpenStudy (anonymous):

no nvm that didn't work sorry

OpenStudy (anonymous):

i was thinking you lost $5 and also the 12% of the $100 for two weeks, but that comes to a loss of $5.50

OpenStudy (mathmate):

Actually, we only lost $95 for two weeks, maybe that would help.

OpenStudy (mathmate):

Actually, approximately, we lost $95 for 1/24th of a year at 0.004733%, which comes to be 0.45, added to $5 we lost in the first gives $5.45, yay!

OpenStudy (mathmate):

The interest rate for (1/24)th of a year is obtained using your formula: r = (1+i/n)^n - 1 with n=24, i=0.12 Check: 1.04733195^24=1.12 (exactly what we want).

OpenStudy (mathmate):

Sorry, it's the other way round: i=(1+r)^(1/n)-1 where i=rate per period, n=24 and your formula should probably read: r = (1+i)^n - 1

OpenStudy (mathmate):

@IStutts i=(1+0.12)^(1/24)-1=0.004733195 Check: 1.004733195^24=1.12000

OpenStudy (anonymous):

ooooh! okay. thank you. sorry, i'm one of those unfortunates that needs a step-by-step.